# arXiv:2601.16212v3[cs.RO] 27 Feb 2026

### v3

@@ 2.2 Data Collection and Generation for Robotics
Robot teleoperation (Mandlekar et al., 2018; Wu et al., 2024; Zhao et al., 2023b; Iyer et al., 2024) is a popular method for collecting task demonstrations – here, humans use a teleoperation device to control a robot and guide it through tasks. Several efforts (Brohan et al., 2022; Ebert et al., 2021; Brohan et al., 2023) have scaled up this paradigm by using a large number of human operators and robot arms over extended periods of time (e.g., months). Some works have also allowed for robot-free data collection with specialized hardware (Chi et al., 2024; Fang et al., 2023; Shafiullah et al., 2023), but human effort is still required for data collection. Other works seek to generate datasets automatically using pre-programmed demonstrators in simulation (Dalal et al., 2023; James et al., 2020; Ha et al., 2023), but scaling these approaches to a larger variety of tasks can be difficult.
@@ 3 Prerequisites
Learning from Demonstrations. The goal of imitation learning is to learn a behavior policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ from a dataset of NN expert demonstrations, denoted as $\mathcal{T}^{e}=\{(o_{t},a_{t})_{t=0}^{T}\}_{n=1}^{N}$, where $o_{t}\in\mathcal{O}$ and $a_{t}\in\mathcal{A}$ represent the observation and action at timestep $t$, and $T$ is the horizon length of each episode. The behavior policy is trained using Behavior Cloning (Pomerleau, 1988) by maximizing the log-likelihood of expert actions, i.e.,
@@ 4.3 Point Extraction
Similar to Haldar and Pinto (2025), we represent the robot end effector as a set of keypoints on the gripper. Given the robot pose $T_{r}^{t}$ at timestep $t$, we define NN rigid transformations $T$ about this pose and compute the pose at each robot keypoint $T_{r}^{t}$ such that

### v4

# Point Bridge: 3D Representations for Cross Domain Policy Learning

@@ 2.2 Data Collection and Generation for Robotics
Acquiring diverse and high-quality demonstrations is a key bottleneck in robot learning. One dominant strategy is teleoperated data collection, in which humans manipulate robots via control devices to perform tasks and record trajectories (Mandlekar et al., 2018; Wu et al., 2024; Zhao et al., 2023b; Iyer et al., 2024). Recent large-scale initiatives (Brohan et al., 2022; Ebert et al., 2021; Brohan et al., 2023) have scaled this approach by coordinating many operators and robots over extended timeframes to build richer datasets. To reduce dependence on physical hardware, alternative systems use purpose-built collection tools that capture demonstrations without a robot directly executing them (Chi et al., 2024; Fang et al., 2023; Shafiullah et al., 2023); however, these pipelines still rely on human input. In contrast, simulation-centric efforts automate data generation with scripted policies or synthetic demonstrators (Dalal et al., 2023; James et al., 2020; Ha et al., 2023), though expanding these methods across varied tasks and real-world conditions continues to be a significant challenge.
@@ 3 Prerequisites
Learning from Demonstrations. The goal of imitation learning is to learn a behavior policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ from a dataset of $N$ expert demonstrations, denoted as $\mathcal{T}^{e}=\{(o_{t},a_{t})_{t=0}^{T}\}_{n=1}^{N}$, where $o_{t}\in\mathcal{O}$ and $a_{t}\in\mathcal{A}$ represent the observation and action at timestep $t$, and $T$ is the horizon length of each episode. The behavior policy is trained using Behavior Cloning (Pomerleau, 1988) by maximizing the log-likelihood of expert actions, i.e.,
@@ 4.3 Point Extraction
Similar to Haldar and Pinto (2025), we represent the robot end effector as a set of keypoints on the gripper. Given the robot pose $T_{r}^{t}$ at timestep $t$, we define $N$ rigid transformations $T$ about this pose and compute the pose at each robot keypoint $T_{r}^{t}$ such that